Let G be a group and H be a non empty finite subset of G.If ab belongs to H for all a,b prove that H is a subgroup of G.will the result be true if H is not finite?
1
Expert's answer
2020-09-24T17:41:58-0400

Recall Subgroup test :

Let "G" be a group and let "H" be a nonempty subset of "G" . If "ab" is in "H"

whenever "a" and "b" are in "H" ("H" is closed under the operation), and "a^{-1}"

is in "H" whenever "a" is in "H" ("H" is closed under taking inverses), then "H"

is a subgroup of "G" .

Solution:

In the view of above test we need only prove that "a^{-1}\in H" whenever

"a\in H" .If "a=e" then "a^{-1}=a" and we are done.

So assume that "a\neq e" . Consider the sequence "a,a^2,a^3,a^4,........." By closer all the elements belong to "H" .

Since "H" is finite , not all of these elements are distinct.

Say "a^i=a^j" where "j>i"

Therefore , "a^{j-i}=e" ; and since "a\neq e" , "j-i>1" .

Thus "aa^{j-i-1}=a^{j-i}=e" .

Therefore,"a^{j-i-1}=a^{-1}" .

But "j-i-1\geq 1" , implies "a^{j-i-1}\in H" and we are done .

If "H" is not finite then result is not true.

Consider the group "G=(\Z,+)" and "H=\N" . Then "a+b\in H , \forall a,b \in H" but "\N" is not a subgroup of "\Z" .